OpenStudy (anonymous):

Elimination??? Select the best possible first step to solving the system by first eliminating the y variable. 4x - 5y = 4 3x + 2y = -20 Multiply the first equation by -2 and multiply the second equation by 5. (i think) Multiply the first equation by 2 and multiply the second equation by -5. Multiply the first equation by 2 and multiply the second equation by 5. None of the above

OpenStudy (anonymous):

X= -4, Y=-4 (-4,-4)

OpenStudy (anonymous):

4x - 5y = 4 (2) 3x + 2y = -20 (5) 8x -10y =8 12 x + 10 y =-100 Can you change your mind now since adding these equations will get rid of the y

OpenStudy (anonymous):

// Solve equation [2] for the variable y [2] 2y = -3x - 20 [2] y = -3x/2 - 10 // Plug this in for variable "y" in equation [1] [1] 4x - 5•(-3x/2-10) = 4 [1] 23x/2 = -46 [1] 23x = -92 // Solve equation [1] for the variable x [1] 23x = - 92 [1] x = - 4 By now we know this much : x = -4 y = -3x/2-10 // Use the x-value to solve for y y = -(3/2)(-4)10 = -4

OpenStudy (anonymous):

Your choice is the third one.

OpenStudy (anonymous):

Hope this helps! :D

OpenStudy (anonymous):

They are not asking to solve the problem. They are asking the first step used to eliminate y